OpenStudy (anonymous):

Solve the system of equations. 4x+3y=13 y=-x+14

OpenStudy (mendicant_bias):

Alright, you can either solve it using substitution or elimination; do you have a preference, or is there a particular way you *have* to do this?

OpenStudy (anonymous):

so you plug in x plus 14 for y and whats 14 times 3

OpenStudy (anonymous):

negative x im sorry

OpenStudy (mendicant_bias):

Going to go with substitution; where \[(1) \ \ \ 4x+3y=13,\]\[(2) \ \ \ y = -x+14,\]Plug in (2) into 1: \[4x+3(-x+14)=13.\]

OpenStudy (anonymous):

@Mendicant_Bias nope

OpenStudy (mendicant_bias):

Alright; do you understand how to solve it from what I've posted above? That last line should be easy to do.

OpenStudy (anonymous):

i think

OpenStudy (mendicant_bias):

Alright, take a shot at it. Get x by itself on one side, and get everything else on the other side.

OpenStudy (anonymous):

cant get it:(

OpenStudy (mendicant_bias):

Do you know how to add and subtract things?

OpenStudy (anonymous):

well yeah

OpenStudy (mendicant_bias):

\[4x+3(-x+14)=13\]

OpenStudy (mendicant_bias):

Alright, when you look at that, what do you think you should first do? First step involves getting rid of those parentheses.

OpenStudy (anonymous):

ok so you distribute the three that gives you -3x and 42 so (4x -3x+ 42 + 13)

OpenStudy (anonymous):

add like terms so then you get 1x + 42=13

OpenStudy (anonymous):

subtract 42 from 13

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

29

OpenStudy (mendicant_bias):

(If you give people the answer, they won't learn.) Glorm did the right thing, I'd answer his question.

OpenStudy (anonymous):

x= -29

OpenStudy (anonymous):

you can check your work by plugging in x